OpenStudy (anonymous):

Helppp, Answer for best answer + Fan Question 1 In simplest form, what is the ratio of triangles to circles? A set of shapes is arranged as follows: triangle, circle, square, circle, circle, square, triangle, circle. one-third one-half seven and two-sevenths one-fourth Question 2 What value of x makes the equation true? two-eighths equals twelve over x ninety-six forty-eight eighteen fourteen Question 3 Which ratio is equivalent to the ratio 8 to 3? two colon one three colon eight sixteen colon six nine co

OpenStudy (anonymous):

two triangles, 4 circles, ratio is \(2:4\) or \(1:2\)

OpenStudy (anonymous):

better known as "one half"

OpenStudy (anonymous):

What is the number question you're answering? Can you include that :) Thanks! But I'm guessing it was #1?

OpenStudy (anonymous):

\[\frac{2}{8}=\frac{12}{x}\] cross multiply, get \[2x=8\times 12\] or \(2x=96\) then divide by 2 and get \[x=96\div 2 = 48\]

OpenStudy (anonymous):

lol yeah, it was #1, best to start at the beginning the next one was #2

OpenStudy (anonymous):

Ah lmao, Np. Thanks

OpenStudy (anonymous):

\[\frac{8}{3}=\frac{8\times 2}{3\times 2}=\frac{16}{6}\]

OpenStudy (anonymous):

Thanks so much! Fanning you + Best response

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Have a good day.